OpenStudy (anonymous):

A vine 6 cm tall is planted in a garden. It grows 4 centimeters a day. How many days will it take to reach 122 cm in height? A. 32 B. 30.5 C. 29 D. 27

OpenStudy (anonymous):

Well, what you would need to do is do 6+4*x. x being each answer choice. So it would be 6+4*32, 6+4*30.5, 6+4*29, 6+4*27. Whichever one equals 122, is the answer.

OpenStudy (anonymous):

let y=height x=number of days y=6+4x for y =122 122=6+4x ==> x=29 so it takes 29 days, the answer is C .

OpenStudy (anonymous):

@ali_rami1990 That's exactly what I got! :)

OpenStudy (anonymous):

Good job then :)

OpenStudy (anonymous):

Hope we helped you, @120